If the mean exam score of a class was 75%, with a standard deviation of 15%, what percent of students would be expected score at or higher than 92%? Assume that the distribution of the scores is normal and the variable is random.
Answer #1

Solution :

Given ,

mean = = 0.75

standard deviation = = 0.15

P(x > 0.92) = 1 - P(x< 0.92)

= 1 - P[(x -) / < (0.92-0.75) / 0.15]

= 1 - P(z <1.13 )

Using z table

= 1 - 0.8708

= 0.1292

answer=12.92%